OpenStudy (anonymous):

The coordinates of rectangle ABCD are A(2, -7), B(8, -7), C(8, -10), D(2, -10). What are the coordinates of the point of intersection of the diagonals of the rectangle after the rectangle is reflected across the x -axis?

OpenStudy (anonymous):

(-5, -8.5) (5, 8.5) (5, -8.5) (-5, 8.5)

OpenStudy (primeralph):

the original diagonal coordinate minus the height

OpenStudy (anonymous):

huh what is the original diagonal ?

OpenStudy (anonymous):

diagonals of a rectangle bisect each other. point of intersection of diagonals=mid point of AC={(2+8)/2 ,(-7-10)/2}=(5,-17/2) Its reflection on X-axis=(5,17/2)

OpenStudy (anonymous):

its reflection on X-axis=(5,8.5)

OpenStudy (anonymous):

Thank you :D